OpenStudy (anonymous):

Help! Limit as x approaches infinity of (sqrt(x^2-x)-x)?

OpenStudy (anonymous):

\[[\lim_{x \rightarrow \infty} \left( \sqrt{x^2-x}-x \right)]\]

OpenStudy (goten77):

-1

OpenStudy (anonymous):

Is this true? and if so, how?

OpenStudy (goten77):

because when you graph it , it shows y is at -1

OpenStudy (anonymous):

Are you able to proof it mathimatically without a graph? That's what I need.

OpenStudy (anonymous):

It is -1/2=-.5

OpenStudy (anonymous):

can someone walk me through the steps?

OpenStudy (anonymous):

\[ \frac {\left( \sqrt{x^2-x}-x \right)\left( \sqrt{x^2-x}+x \right)} {\left( \sqrt{x^2-x}+x \right)}=\\ \frac{x^2-x-x^2}{\left( \sqrt{x^2-x}+x \right)}=\frac {-x}{\left( \sqrt{x^2-x}+x \right)} \] The last expression goes to -1/2 when x goes to Infinity

OpenStudy (anonymous):

ok up to here. I understand now. Now, how did you get -1/2 when it goes to infinity?

OpenStudy (anonymous):

\[ \left( \sqrt{x^2-x}-x \right)=\frac {\left( \sqrt{x^2-x}-x \right)\left( \sqrt{x^2-x}+x \right)} {\left( \sqrt{x^2-x}+x \right)}=\\ \frac{x^2-x-x^2}{\left( \sqrt{x^2-x}+x \right)}=\frac {-x}{\left( \sqrt{x^2-x}+x \right)}=\\ \frac {-1}{\left( \sqrt{1-\frac 1 x}+1 \right)} \]

OpenStudy (anonymous):

For the last equality divide up and down by x

OpenStudy (goten77):

this guy is 2 smart

OpenStudy (anonymous):

\[ \frac {-1}{\left( \sqrt{1-\frac 1 x}+1 \right)} \] goes to -1/2 when x goes to Infinity

OpenStudy (anonymous):

Did you get it?

OpenStudy (goten77):

wait eliassaab when you did that multiplying did you multiply by the conjuant? or however its spelled

OpenStudy (anonymous):

yes, I think so. IN other words when the -1/x goes to infinity, it is such a negligible difference that we don't count it. It all makes sense now. You are a life savor!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you know how to solve functions that have a floor function in them as well?

OpenStudy (anonymous):

Close this problem and ask the new one with floor function.

OpenStudy (anonymous):

ok